OpenStudy (anonymous):

If Sachi, Sheryl, Sharon, Shannon, and Stephanie are all sitting in the back seat of the school bus, in how many different ways can they seat themselves? Is it just one seat?

OpenStudy (anonymous):

But there are 5 girls.... so 5*4*3*2*1=120. But that would be the same if there was 2 seats or 4 seats right?

OpenStudy (anonymous):

Hmm, I'm sorry, there are five girls. Then what you do is: Well, assuming there are five seats in the back of the bus. Each girl must have one seat, and cannot have more than one. So what you do is multiply the options together! So the first seat can have 1 out of the five of them. So the second seat can have 1 out of the four of them. (remember one is already sitting) So the third seat can have 1 out of the three of them. (so two were sitting before) So the fourth seat can have 1 out of the two of them. (so three were sitting before) And the final seat has the last one. So you multiply the number of choices together for the seats to get: 5*4*3*2 = 120 So there are 120 possible arrangements.

OpenStudy (anonymous):

No, it would only work for 5 seats or 4 seats., If there were two seats, only 2 girls could sit right? So it becomes 5 girls (for first seat) * 4 girls (for second seat) = 20 choices.

OpenStudy (anonymous):

I get what your saying, but I thought the question was asking how many combinations if they were sitting together. Hmmmm,

OpenStudy (anonymous):

No, there are a 120 combinations like I showed you since they're all sitting. You have 5 seats.